A 200-turn rectangular coil having dimensions of 3.0cm by 6.0cm is placed in a uniform magnetic field of magnitude 0.89T. Find the current in the coil if the maximum torque exerted on it by the magnetic field is 0.14N*m
Answer #1

=

Torque maximum for theta=90

Angle between magnetic moment and area vector=90

m=current × area =nIA=200×I×3.0cm×6cm =current×0.36

B=.89T

Torque(maximum)=m×B=.14

Current×0.3204=.14

I=0.14/0.3204=0.437 amper